Suppose that license tags in another state are made up of three letters followed by three digits (0 - 9), none of which can repeat. Additionally, each tag number may be assigned to a regular tag, a wildlife-conservation tag, or a veterans tag.
(a) How many unique tags can be made?
(b) What’s the probability that you are parked next to a car from this state that has a tag that ends in a 9?
Answer #1

a) number of ways =N(choose 3 letter from 26 and arrange, choose 3 digits from 10 and arrange and choose 1 tag from 3) =26*25*24*10*9*8*3 =33696000

b)

since last digit has 10 choices ,

therefore

probability that you are parked next to a car from this state that has a tag that ends in a 9 =1/10 =0.1